OpenStudy (anonymous):

Carter has 37 coins, all nickels and dimes in the piggy bank. The value of the coins is $3.10. How many dimes does carter have?

OpenStudy (freckles):

so one equation can be made from the first sentence and another equation can be made from the second sentence let N= number of nickels let D= number of dimes you are given N+D is?

OpenStudy (freckles):

how many coins does carter have?

OpenStudy (anonymous):

37 so do you put D+N=37?

OpenStudy (freckles):

yes because that is all he has is dimes and nickels

OpenStudy (freckles):

now a nickel is 5 cents or .05 dollars and a dime is 10 cents or .10 dollars also can be read as .1 dollars

OpenStudy (freckles):

so we are given .05*N+.1*D=?

OpenStudy (anonymous):

.05n+.1d=3.10

OpenStudy (freckles):

right \[\text{ we have the following system \to solve } \\ N+D=37 \\ .05 N+.1D=3.1\]

OpenStudy (freckles):

if you don't like the decimal action in the second equation multiply both sides by 100 giving you 5N+10D=310

OpenStudy (freckles):

you could also divide both sides of that equation by 5 giving you N+2D=310/5

OpenStudy (freckles):

310/5=62

OpenStudy (freckles):

you have the system is: N+D=37 N+2D=62 surely this looks less ugly to solve

OpenStudy (freckles):

the question ask you to solve for the number of dimes aka what we called D

OpenStudy (freckles):

You could solve the first equation for N and then replace the second equation's N with what you found from the first

OpenStudy (anonymous):

ok i got it mow thanks for the help

OpenStudy (freckles):

meow

OpenStudy (freckles):

i mean np

OpenStudy (freckles):

sorry mow made me think of meow i know you meant now though